a body of mass 0.004kg executes SHM the force acting on the body when the displacement is 8cm is 0.024kgwt find the periodif the maximum velocity is 5m s find amplitude and maximum acceleration given g 9.8 m s²

1
Expert's answer
2022-04-11T12:07:26-0400

"k=\frac Fx=0.3~\frac N{kg},"

"\omega=\sqrt\frac km=8.7~\frac{rad}s,"

"T=\frac {2\pi}{\omega}=0.73~s,"

"A=\frac v{\omega}=0.57~m,"

"a=\omega v=43~\frac m{s^2}."